what is the x-intercept of y=1/3x -6

Answer by Cromlix(4381)

To find 'x' intercept make y = 0
y = 1/3x - 6
0 = 1/3x - 6
1/3x = 6
x = 18
So, 'x' intercept = 18,0
Hope this helps.
:-)